# Binary Autoencoder for Mechanistic Interpretability of Large Language Models

### v1

@@ Appendix B Interpretability Score Distribution: Selecting a Better Interpretable Feature Set
Llama 3.2-1B Layer 11 $d=2048$
ReLU SAE 2065 1380 0.260
@@
Llama 3.2-1B Layer 14 $d=2048$
ReLU SAE 2528 1600 0.217
@@
Llama 3.2-3B Layer 20 $d=3072$
ReLU SAE 1923 1289 0.312
@@
Appendix C Case Analysis: Extracted Features
@@
Feature index: #2289 Feature interpretation: “verbs of visual attention or perception” Interpretability score: 0.875 Activate tokens & contexts: #1: “but just one look at a minion sent him practically catatonic .” #2: “each time she looked into mason ’s face , she was grateful that he looked nothing like his father .” #3: “megan asked , gazing from noah ’s…C.2 #4: “peeking out from the covering , she saw emma was wearing her signature color , green .” #5: “of course , he ’d probably argue that while the gown might not have held up , he still looks fabulous and much younger than his age .” #6: “megan glanced between the two of them .” #7: “she stared into his face before she responded .”
@@
C.1 High-score Features
@@
C.2 Low-score Features
1010footnotetext: To comply with the requirements of the ethics review, we have redacted potentially offensive or NSFW content from these input cases.
Also, we list some features with low interpretability scores, as shown in Fig. 1619. Among these low-score features, we summarize the characteristics of these features: (1) Subwords from the tokenizer. As shown in Fig. 1719, these subword tokens share no obvious commonality even for humans, and they greatly confuse the LLM, leading it to produce irrelevant interpretations. This illustrates the drawback of token-based interpretation, as noted in our limitations. (2) Function words with similar context. As shown in Fig. 1618, these function-word tokens, despite appearing in similar contextual environments (often originating from the same sentence or even the same region within a sentence), lack surface-level commonality and are therefore difficult to interpret by ComSem. This highlights the urgency of interpreting directly from the decoded feature vectors instead of original tokens. Moreover, from this phenomenon, a further hypothesis is that function words without clear semantics may be more likely to concentrate contextual information during contextualization. (3) Character level similarity. As shown in Fig. 19, the tokens activated on feature 5949 all end with “em”, according to previous works (Zhang & He, 2024; Fu et al., 2024), backend LLM may struggle to process such inputs, causing a low interpretability score.
C.3 Scale of High-score Features and Low-score Features
@@
In detail, we conduct UMAP (McInnes et al., 2018), a dimensionality reduction based on the macroscopic adjacency structure from Euclidean distance, on hidden states of all the tokens from the sentences whose token activates the investigated feature, as shown in Fig. 40 for 3 high-score features, and Fig. 41 for 3 low-score features, where an important observation is that: the hidden states activated by high-scoring features form clear clusters, suggesting macroscopic spatial similarity. In contrast, the hidden states activated by low-scoring features are more dispersed overall, exhibiting similarity only within certain subspaces (i.e., along the corresponding rows of the activated features in $W_{\textrm{out}}$). That is, the low-score features are more folded and implicit, so that harder to extract.
Appendix D Augmentation Experiments and Results
D.1 Activation Magnitude Visualization of BAE and SAE
@@
D.3 More Cases for Fig. 6
We extend the investigation scope of Fig. 6, where the activation frequency of each feature is visualized, as shown in Fig. $39$. The conclusion derived from these augmentation results is consistent with the main body.
@@
11    $\mathcal{D}_{I}[i]\coloneqq\{\texttt{"Interp" = }\mathrm{LM}(\mathrm{Prompt}_{\mathrm{Interp}}(\mathcal{D}_{S}[i][1:n_{I}])),\texttt{ "Activated" = True}\}$
12    $\mathrm{FA}\coloneqq\mathrm{FA}+1$
13 else
@@ 12    $\mathrm{FA}\coloneqq\mathrm{FA}+1$
14    $\mathcal{D}_{I}[i]\coloneqq\{\texttt{"Interp" = None, "Activated" = False}\}$
15   end if
16
17 end for
18
@@ 14    $\mathcal{D}_{I}[i]\coloneqq\{\texttt{"Interp" = None, "Activated" = False}\}$
19 for $i\coloneqq 1$ to $d^{\prime}$ do
20 if $\mathcal{D}_{I}[i][\texttt{"Activated"}]$ then
21    $\mathcal{D}_{I}[i][\texttt{"Score"}]\coloneqq 0$
22    for $(t^{(j)},c^{(j)})$ in $\mathcal{D}_{S}[i][n_{I}+1:n_{I}+n_{T}]$ do
@@ 21    $\mathcal{D}_{I}[i][\texttt{"Score"}]\coloneqq 0$
23       if $\mathrm{LM}(\mathrm{Prompt}_{\mathrm{Test}}((t^{(j)},c^{(j)}),\mathcal{D}_{I}[i][\texttt{"Interp"}]$) then
24          $\mathcal{D}_{I}[i][\texttt{"Score"}]=\mathcal{D}_{I}[i][\texttt{"Score"}]+1$
25         end if
26
27      end for
28    $\mathcal{D}_{I}[i][\texttt{"Score"}]=\mathcal{D}_{I}[i][\texttt{"Score"}]/\texttt{len(}\mathcal{D}_{S}[i][n_{I}+1:]\texttt{)}$
29    $\mathrm{Score}\coloneqq\mathrm{Score}+\mathcal{D}_{I}[i][\texttt{"Score"}]$
30    if $\mathcal{D}_{I}[i][\texttt{"Score"}]>0$ then
31
32       $\mathrm{FI}\coloneqq\mathrm{FI}+1$
33      end if
34
35   end if
36
37 end for

### v2

@@ Appendix D Natural Low-entropy Tendency of BAE Training
Llama 3.2-1B Layer 11 $d=2048$ ReLU SAE 2065 1380 0.260
@@
Llama 3.2-1B Layer 14 $d=2048$ ReLU SAE 2528 1600 0.217
@@
Llama 3.2-3B Layer 20 $d=3072$ ReLU SAE 1923 1289 0.312
@@
Moreover, interestingly, when the estimated entropy constraint is disabled (red in Fig. 2), the entropy gathers higher but near the diagonal, suggesting that BAE without entropy training objective can find a relatively low-entropy encoding $h_{B}$, even if no normalizations are conducted to minimize the encoding entropy. To get a closer observation, we plot the self-regression loss $\mathcal{L}_{r}$ and training-time entropy $\mathcal{L}_{e}$ on every training step, with $r=2$ (more cases in Appendix G.2), as shown in Fig. 12, with normal setting (left) or $\alpha_{e}=\alpha_{c}=0$ (right). In the normal setting (left), a stable and monotonous convergence to both minimal entropy and self-regression loss can be observed. However, if the entropy loss is disabled (right), the training dynamics become multiphase and non-monotonic. Specifically, Phase 1: the loss rapidly decreases while the hidden activations maintain relatively high entropy; Phase 2: entropy sharply drops to near zero, with the loss remaining low, even if no entropy and weight penalty are utilized; and Phase 3: both loss and entropy oscillate in a narrow range at the end of training. Such an observation suggests that: Gradient descent on simple regression loss finds relatively “simple” representations, after a long-term stagnation on loss value, supporting previous work on training dynamics (tishby2015deep; michael2018on; huh2023the; nanda2023progress; shah2020pitfalls; frankle2018the; bartlett2020benign). Additionally, enabling the entropy penalty suppresses the harmful Phase 3 oscillations, confirming its effectiveness.
Appendix E Proof: More Burst Features Carry More Information
This section substantiates our claim that features with higher burstiness contain more information and should therefore be preferentially selected as activated features. Our proof is grounded in fundamental principles of information theory:
Theorem E.1 (Burst Features Carry More Information).
Let $X_{1}\sim\mathrm{Bernoulli}(p_{1})$, $X_{2}\sim\mathrm{Bernoulli}(p_{2})$, the information of an observation (i.e., the actual value of hidden activation in a specified channel) $x\in\{0,1\}$ can be written as $I_{X}(x):=-\log\Pr[X=x]$. Then:
$I_{X_{1}}(x_{1})\geqslant I_{X_{2}}(x_{2})\Leftrightarrow|x_{1}-p_{1}|\geqslant|x_{2}-p_{2}|$
The proof is straightforward. That is, a more bursty channel carries more information, while our estimator of the total entropy from the hidden state set is fixed. Therefore, identifying these more bursty features is more conducive to reconstructing this particular observation from the average (i.e., the typical set).
Appendix F Case Analysis: Extracted Features
@@
Feature index: #2289 Feature interpretation: “verbs of visual attention or perception” Interpretability score: 0.875 Activate tokens & contexts: #1: “but just one look at a minion sent him practically catatonic .” #2: “each time she looked into mason ’s face , she was grateful that he looked nothing like his father .” #3: “megan asked , gazing from noah ’s…F.2 #4: “peeking out from the covering , she saw emma was wearing her signature color , green .” #5: “of course , he ’d probably argue that while the gown might not have held up , he still looks fabulous and much younger than his age .” #6: “megan glanced between the two of them .” #7: “she stared into his face before she responded .”
@@
F.1 High-score Features
@@
F.2 Low-score Features
99footnotetext: To comply with the requirements of the ethics review, we have redacted potentially offensive or NSFW content from these input cases.
Also, we list some features with low interpretability scores, as shown in Fig. 1821. Among these low-score features, we summarize the characteristics of these features: (1) Subwords from the tokenizer. As shown in Fig. 1921, these subword tokens share no obvious commonality even for humans, and they greatly confuse the LLM, leading it to produce irrelevant interpretations. This illustrates the drawback of token-based interpretation, as noted in our limitations. (2) Function words with similar context. As shown in Fig. 1820, these function-word tokens, despite appearing in similar contextual environments (often originating from the same sentence or even the same region within a sentence), lack surface-level commonality and are therefore difficult to interpret by ComSem. This highlights the urgency of interpreting directly from the decoded feature vectors instead of original tokens. Moreover, from this phenomenon, a further hypothesis is that function words without clear semantics may be more likely to concentrate contextual information during contextualization. (3) Character level similarity. As shown in Fig. 21, the tokens activated on feature 5949 all end with “em”, according to previous works (zhang-he-2024-large; fu2024large), backend LLM may struggle to process such inputs, causing a low interpretability score.
F.3 Scale of High-score Features and Low-score Features
@@
In detail, we conduct UMAP (mcinnes2018umap), a dimensionality reduction based on the macroscopic adjacency structure from Euclidean distance, on hidden states of all the tokens from the sentences whose token activates the investigated feature, as shown in Fig. 42 for 3 high-score features, and Fig. 43 for 3 low-score features, where an important observation is that: the hidden states activated by high-scoring features form clear clusters, suggesting macroscopic spatial similarity. In contrast, the hidden states activated by low-scoring features are more dispersed overall, exhibiting similarity only within certain subspaces (i.e., along the corresponding rows of the activated features in $W_{\textrm{out}}$). That is, the low-score features are more folded and implicit, so that harder to extract.
Appendix G Augmentation Experiments and Results
G.1 Activation Magnitude Visualization of BAE and SAE
@@
G.3 More Cases for Fig. 5
We extend the investigation scope of Fig. 5, where the activation frequency of each feature is visualized, as shown in Fig. $41$. The conclusion derived from these augmentation results is consistent with the main body.
@@
     $\mathcal{D}_{I}[i]\coloneqq\{\texttt{"Interp" = }\mathrm{LM}(\mathrm{Prompt}_{\mathrm{Interp}}(\mathcal{D}_{S}[i][1:n_{I}])),\texttt{ "Activated" = True}\}$
    $\mathrm{FA}\coloneqq\mathrm{FA}+1$
 else
@@
     $\mathcal{D}_{I}[i]\coloneqq\{\texttt{"Interp" = None, "Activated" = False}\}$
  end if
end for
@@
for $i\coloneqq 1$ to $d^{\prime}$ do
  if $\mathcal{D}_{I}[i][\texttt{"Activated"}]$ then
     $\mathcal{D}_{I}[i][\texttt{"Score"}]\coloneqq 0$
    for $(t^{(j)},c^{(j)})$ in $\mathcal{D}_{S}[i][n_{I}+1:n_{I}+n_{T}]$ do
@@
        if $\mathrm{LM}(\mathrm{Prompt}_{\mathrm{Test}}((t^{(j)},c^{(j)}),\mathcal{D}_{I}[i][\texttt{"Interp"}]$) then
          $\mathcal{D}_{I}[i][\texttt{"Score"}]=\mathcal{D}_{I}[i][\texttt{"Score"}]+1$
        end if
     end for
    $\mathcal{D}_{I}[i][\texttt{"Score"}]=\mathcal{D}_{I}[i][\texttt{"Score"}]/\texttt{len(}\mathcal{D}_{S}[i][n_{I}+1:]\texttt{)}$
    $\mathrm{Score}\coloneqq\mathrm{Score}+\mathcal{D}_{I}[i][\texttt{"Score"}]$
    if $\mathcal{D}_{I}[i][\texttt{"Score"}]>0$ then
       $\mathrm{FI}\coloneqq\mathrm{FI}+1$
     end if
  end if
end for